OpenStudy (anonymous):

Systems of equatons: Solve for x and y y=2x+9 y=x-5

OpenStudy (solomonzelman):

multiply the second equation times -2 or -1 and add the equations

OpenStudy (solomonzelman):

when subtracting equations might be confusing, multiply by negative and add, to avoid confusion.

OpenStudy (solomonzelman):

or you can use a matrix, which would be fun too.

OpenStudy (anonymous):

Jeez I am only in 8th grade. This is freaking Confusing xD

OpenStudy (anonymous):

@SolomonZelman So when i multiply the 2nd one i got -2x +10. Is that correct?

OpenStudy (solomonzelman):

when you multiply the second equation times -2 , you have: \(\large\color{slate}{\displaystyle y=x-5}\) \(\large\color{slate}{\displaystyle \left[y\times -2\right]=\left[x\times -2\right]\left[-5\times -2\right]}\) \(\large\color{slate}{\displaystyle -2y=-2x +10}\) correct!

OpenStudy (solomonzelman):

Now, add the equations together: \(\large\color{slate}{\displaystyle ~~~-2y~=~-2x~~+10}\) \(\large\color{slate}{\displaystyle ^{^{\LARGE+}}~~~~~~y~~=~~~2x~~~~~+9}\) \(\Large\color{slate}{\displaystyle ^{\text{______________________}}}\)

OpenStudy (solomonzelman):

do each thing separately. 1. Add the y's 2. Add the x's 3. Add the numbers and re-write the equation.

OpenStudy (anonymous):

y= 19 did i do that right?

OpenStudy (solomonzelman):

almost.

OpenStudy (anonymous):

Wait its -y = 19

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

and if -y=19 then y=? (can you tell me that?)

OpenStudy (solomonzelman):

(multiply both sides times -1 to find what y is)

OpenStudy (anonymous):

-19?

OpenStudy (solomonzelman):

yes y=-19

OpenStudy (anonymous):

okay now I need to find x

OpenStudy (solomonzelman):

Now, after you found y, you can solve for x without a problem by plugging your y-value (which is -19) into ANY of the ORIGINAL EQUATION.

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle y=2x+9}\) \(\large\color{slate}{\displaystyle \color{red}{-19}=2x+9}\) and then sovle for x.

OpenStudy (solomonzelman):

solve*

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle 19=2x+9}\) \(\large\color{slate}{\displaystyle 19\color{red}{-9}=2x+9\color{red}{-9}}\) \(\large\color{slate}{\displaystyle 10=2x}\) and then, \(\large\color{slate}{\displaystyle x=?}\)

OpenStudy (anonymous):

x= 5

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Thank you @SolomonZelman Can you help me with one more?

OpenStudy (solomonzelman):

So your solution is x=5 y=-19 In other words, your 2 equations y=2x+9 y=x-5 intersect at a point (5,-19)

OpenStudy (solomonzelman):

yes, I can help with another one.

OpenStudy (anonymous):

y= x - 2 2x - 3y = -5

OpenStudy (solomonzelman):

yes. The way they are writing the system, they are wanting you to substitute (do 'substitution').

OpenStudy (solomonzelman):

Your first equation \(\large\color{black}{ \displaystyle y= x - 2 }\), is telling us, that (in this case) \(\large\color{black}{ \displaystyle x - 2 }\) is the same exact thing as \(\large\color{black}{ \displaystyle y }\). That means wherever you see a \(\large\color{black}{ \displaystyle y }\) in the other equation, you can replace this \(\large\color{black}{ \displaystyle y }\) by \(\large\color{black}{ \displaystyle x - 2 }\).

OpenStudy (solomonzelman):

The way this will happen is as follows: your second equation. \(\large\color{black}{ \displaystyle 2x - 3y = -5 }\) replacing the ` y ` with ` x+2 ` \(\large\color{black}{ \displaystyle 2x - 3(\color{blue}{x+2}) = -5 }\)

OpenStudy (solomonzelman):

See what is going on here up to this point?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Wait so x = -1

OpenStudy (anonymous):

@SolomonZelman are all of these in the right quads? http://gyazo.com/d89273d94b017f8d6a9f0a1fe739ffb4

OpenStudy (solomonzelman):

you can graph the equations on https://www.desmos.com/calculator and tell me in which quadrant the intersection point occurs.

OpenStudy (solomonzelman):

is this graphing calculator good ?

OpenStudy (anonymous):

2x-y=-2 and x - 2y = 5 intercept at points (-3,-4)